Dilly Farm Supply is located in a small town in the rural west. Data regarding the store's operations follow:

  • Sales are budgeted at $290,000 for November, $310,000 for December, and $210,000 for January.
  • Collections are expected to be 65% in the month of sale and 35% in the month following the sale.
  • The cost of goods sold is 80% of sales.
  • The company desires to have an ending merchandise inventory at the end of each month equal to 70% of the next month's cost of goods sold. Payment for merchandise is made in the month following the purchase.
  • Other monthly expenses to be paid in cash are $21,100.
  • Monthly depreciation is $21,000.
  • Ignore taxes.
Balance Sheet
October 31
Assets
Cash $ 25,000
Accounts receivable 77,000
Merchandise inventory 162,400
Property, plant and equipment, net of $624,000 accumulated depreciation 1,026,000
Total assets $ 1,290,400
Liabilities and Stockholders' Equity
Accounts payable $ 239,000
Common stock 740,000
Retained earnings 311,400
Total liabilities and stockholders' equity $ 1,290,400

The difference between cash receipts and cash disbursements for December would be:

Multiple Choice

  • $38,700

  • $13,700

  • $46,600

  • $19,200
Answer #1

Cash receipt = (290000*35%+310000*65%) = 303000

cash disbursement

Purchase November = (290000*80%)+(310000*80%*70%)-162400 = 243200

Total cash disbursement = 243200+21100 = 264300

Difference between Cash receipts and cash disbursement = 303000-264300 = 38700

So answer is a) $38700